OpenStudy (anonymous):

Solve for y in the following system of equations. 6x - 4y = -12 -6x + 2y = 0 i really need help anyone can you help me

OpenStudy (anonymous):

add the two equations and post what you get

OpenStudy (anonymous):

see i dont even no how to do that. and i really need help

OpenStudy (anonymous):

if a = 2 and b = 3, then a+b = 2+3 a+b = 5 similarly add the two equations.

OpenStudy (anonymous):

do you see the logic of adding the equations? now try it on your own and post what you get.

OpenStudy (anonymous):

x=2 y=6

OpenStudy (anonymous):

He means to combine the two, add everything on the left of the two equations together, and set all of that = to everything on the right of both equations added together

OpenStudy (anonymous):

if u havent learnt how to do that yet u can solve for y in the first equation and replace y in the second equation: 1st equation : 6x-4y=-12 solve for y: y=3/2x+3 Now replace the y in the 2nd equation (-6x+2y=0) so it becomes:-6x+2(3/2x+3)=0 and solve from there

OpenStudy (anonymous):

6x-4y=-12 -6x+2y=0 adding them,we get y=6

OpenStudy (anonymous):

what destinc and dhatraditya want u to do is add the equations together. so u take the 1st equation 6x-4y=-12 Add to the left side of the equation -6x+2y so it becomes: 6x-4y+(-6x+2y) which equals: -2y now when u add something to the left side of an equation u have to add it to the right side as well so on the right side u should add -12+(-6x+2y) however from the 2nd equation u know that -6x+2y=0 so u can replace the"-6x+2y" on the right side with 0 because the 2nd equations says they are equal so it becomes -12+0 so in the end u wind up with: -2y=-12